Given the matrices (Each matrices are inside a [ ] )
3 0 1 5 2 0 1 3
A= -1 2 , B= 4 -1 , C= 1 4 2 , D= -1 0 1 , E= -1 1 2
1 1 0 2 3 1 5 3 2 4 4 1 3

Compute if possible
1. -4C^2
2. (E-D)^T
3. A(BC)
4. (B^T B)C
5. 3B-AB
1
Expert's answer
2020-05-29T16:46:25-0400

"A=\begin{pmatrix}
 3 & 0 \\
 -1 & 2\\
1&1
\end{pmatrix}, B=\begin{pmatrix}
 1 & 5 \\
 4 & -1\\
0&2
\end{pmatrix}, \\
C=\begin{pmatrix}
 2 & 0&1 \\
 1 & 4&2\\
3&1&5
\end{pmatrix}, D=\begin{pmatrix}
 -1 &0&1 \\
 3 & 2&4
\end{pmatrix}, \\
E=\begin{pmatrix}
 -1 &1&2 \\
 4 & 1&3
\end{pmatrix}"

"-4C^2=-4\begin{pmatrix}
 2 & 0&1 \\
 1 & 4&2\\
3&1&5
\end{pmatrix}\cdot\begin{pmatrix}
 2 & 0&1 \\
 1 & 4&2\\
3&1&5
\end{pmatrix}=\\
=-4\begin{pmatrix}
 4+0+3 & 0+0+1&2+0+5 \\
 2+4+6 & 0+16+2&1+8+10\\
6+1+15&0+4+5&3+2+25
\end{pmatrix}=\\
=-4\begin{pmatrix}
 7 & 1&7 \\
 12 & 18&19\\
22&9&30
\end{pmatrix}=\\
=\begin{pmatrix}
 - 28 & -4&-28 \\
 -48 & -72&-76\\
-88&-36&-120
\end{pmatrix}"

2.

"(E-D)^T=\\=\left(\begin{pmatrix}
 -1 &1&2 \\
 4 & 1&3
\end{pmatrix}-\begin{pmatrix}
 -1 &0&1 \\
 3 & 2&4
\end{pmatrix}\right)^T=\\
=\left(\begin{pmatrix}
 0 &1&1 \\
 1 & -1&-1
\end{pmatrix}\right)^T=\\
=\begin{pmatrix}
 0 & 1 \\
 1& -1\\
1&-1
\end{pmatrix}"


3.

"A(BC)"

"B:3\times2,
C:3\times3\\"

The product "BC" does not exist.

4.

"(B^TB)C"

"B^T=\begin{pmatrix}
 1 & 4&0 \\
 5 & -1&2
\end{pmatrix},B=\begin{pmatrix}
 1 & 5 \\
 4 & -1\\
0&2
\end{pmatrix}\\
B^T\cdot B=\begin{pmatrix}
 1 & 4&0 \\
 5 & -1&2
\end{pmatrix}\cdot\begin{pmatrix}
 1 & 5 \\
 4 & -1\\
0&2
\end{pmatrix}=\\
=
\begin{pmatrix}
 1+16+0 & 5-4+0 \\
 5-4+0 & 25+1+4
\end{pmatrix}=\begin{pmatrix}
 17 & 1 \\
 1 & 30
\end{pmatrix}"

"B^T\cdot B:2\times2, C:3\times3"

The product "(B^TB)C" does not exist.

5.

"3B-AB\\
3B=3\begin{pmatrix}
 1 & 5 \\
 4 & -1\\
0&2
\end{pmatrix}=\begin{pmatrix}
 3 & 15 \\
 12 & -3\\
0&6
\end{pmatrix}"

"A:3\times2, B:3\times2"

The product "AB" does not exist.

Then 3B-AB does not exist.